Please show all steps
(c) Hence or otherwise solve the equation use vat substitution
Answer #1

2 take y = y ² 3. - - - у + (-5) = (12) 22 P(X) = -1 910) - factor is. → integrating px) dx 4x)=e - 1 dx =e - -enx general solution is 1.9() +C у ) (© аз к ucal, -1 ་) ༧༣ ,,, sk ___ ༤ ན 、, | , ན V- , +cམ put back v=y² ༤- ད༔ t Ita2 ་ | +cxt